a company manufactures a television which is found to have a 10% probability of being defective in some way. a sample of 4 televisions is taken.
1) what is the probability that exactly 1 television will be defective?
2) what is the mean of this distribution?
3) what is the standard deviation of this distribution?
having issues with these three problems, thank you

Answer by stanbon(75887)

a company manufactures a television which is found to have a 10% probability of being defective in some way. a sample of 4 televisions is taken.
---
Binomial Problem with n = 4 and p = 0.1
------
1) what is the probability that exactly 1 television will be defective?
P(x=1) = 4C1(0.1)*(0.9)^3 = 0.2916
-------------------------------
2) what is the mean of this distribution?
u = np = 4*0.1 = 0.4
-------------------------------
3) what is the standard deviation of this distribution?
std = sqrt[npq] = sqrt[0.4*0.9] = 0.6
===========================================
Cheers,
Stan H.
=============